OpenStudy (anonymous):

I need to graph the inequality of a plane for: 2x+3y≤6

OpenStudy (watchmath):

First you solve for \(y\) \(y\leq -\frac{2}{3}x+2\) Then draw the line \(y=-\frac{2}{3}x+2\) The region described by the inequality is exactly the region below the line.